OpenStudy (jesuslover11):

Solve the equation. 5x+2=62 a)11.5 b)12 c)12.8 d)36

OpenStudy (anonymous):

5x+2=62 5x=62-2 5x=60 x=\[\frac{ 60 }{ 5 }\] x= 12

OpenStudy (jesuslover11):

thank you i have another though

OpenStudy (anonymous):

the answer is b) 12

OpenStudy (jesuslover11):

x/4+1=-5 a) -2 b)16 c)4 d)-24

OpenStudy (jesuslover11):

help

OpenStudy (anonymous):

-24

OpenStudy (jesuslover11):

thnk you can you do another for me

OpenStudy (anonymous):

Sure

OpenStudy (jesuslover11):

complete the function table. Then write a rule for the function.